OpenStudy (anonymous):

Use the method of subsitution to solve the sysytem of linear equations. 2x-y=3 7x-3y=13

OpenStudy (anonymous):

2x-y = 3 => y = 2x-3 Now, substitute this value of y in the second equation, and solve it - 7x - 3(2x-3) = 13

OpenStudy (hba):

@Baldwin777 Start solving this and get x

OpenStudy (anonymous):

okay so I would get 7x-6x-3=13 right?

OpenStudy (hba):

No @Baldwin777 You need to know something (-)(-)=+ (+)(-)=- (+)(+)=+ So now i would recommend to do it again 7x - 3(2x-3) = 13 I am here so i can help :)

OpenStudy (anonymous):

No, You tried well though! How much is -3 x -3 ?

OpenStudy (hba):